Alpha Assemblies produces a wide variety of products in their central Nebraska facility. Many of these products fit within a line of standard truck and trailer parts. Within this is a set of hitches that can be made in two sizes (1-1/4 and 2) and three strengths (light duty, medium duty, and heavy duty). All of the hitches in this line have "natural" finishes so no additional steps are necessary after they are cast. The machine used to cast these six hitch variations is available for 2000 hours per year. Operators are paid $15.00 per hour including benefits.

Simone is working to establish her reputation at Alpha Assemblies as part of her internship opportunity. She is looking for a way to improve the operations. She has been working on the hitch project and is looking at possible ways to cut costs. At this point she has studied the processes and their associated costs, and completed forecasts for the next year's demand.

Hitch A is the 1-1/4 inch, light duty, natural hitch. Simone has forecast the demand for this hitch at 6,000 units for the next year. It takes one operator 5.7 hours to perform a setup to produce a batch for this hitch and no other setup costs are incurred when setting up the machine. In addition, the holding costs for this item are $4.16 per unit per year. This hitch has a current batch size of 1,500 units per production run.

Hitch B is the 1-1/4 inch, medium duty, natural hitch. This hitch has a forecast annual demand for next year of 3,000 units and has a current batch size of 1,500 units per production run. The unit processing time for this hitch is 0.0420 hours per unit, setup time using a single operator is 6.6 hours per setup and the holding cost per unit per year is $3.34. No additional costs are incurred in the setup process.

Hitch C has a forecast annual demand for next year of 7,000 units and a current batch size of 1,500 units. It takes 0.0400 hours to produce each unit and the setup time for a production run is 8.4 hours. Hitch C has a holding cost of $3.57 per unit per year. The value of each Hitch C as it is stored in inventory is $29.66 per unit. As with Hitches A and B, it takes 1 worker to perform the setup and no other costs are incurred in the setup process.

Hitch D has a forecast annual demand for the next year of 10,000 units. It takes 5.7 hours to complete a setup using a single worker and there are no other costs incurred during the setup process. The holding cost per unit per year is $3.57. The current batch size is 2,000 units. The unit processing time for Hitch D is 0.0380 hours per unit.

How much more time (in hours) will be spent on setups if Hitch D is produced using Simone's recommendation of EOQ than with the current batch size?
Answer #1

No. of batches needed to complete the demand = 1000/2000 = 5

Time taken to produce one batch = 5.7 + 2000 * 0.0380 hours = 81.7 hours

Total time taken = 5*81.7 = 408.5

Labor Cost = $15*408.5 = $6127.5

Holding cost = $3.57

Fixed Cost = 6127.5 + 3.57 = $6131.07

EOQ = Sqrt. (2FD/C) = Sqrt.([2*6131.07*10000]/15) = 2859